A microwave oven cooks food with microwaves of wavelength 11.8 cm. What is the frequency of these microwaves?
Answer #1

We know that \(c=f \lambda\)

We have \(c=\) speed of light \(=3 \times 10^{8} \mathrm{~m} / \mathrm{sec}\)

\(\lambda=11.8 \mathrm{~cm}=11.8 \times 10^{-2} \mathrm{~m}\)

Thus \(f=c / \lambda\)

\(f=3 \times 10^{8} / 11.8 \times 10^{-2} \mathrm{~s}^{-1}\)

\(\mathrm{f}=2.54 \times 10^{9} \mathrm{~Hz}\)